# Two Causally Related Needles in a Video Haystack

### v1

@@ 4.4 Analysis of Dataset Bias
To prevent performance underestimation due to the out-of-distribution visual grounding format, our dataset introduces an image description (ID) format. However, a VLM may have memorized details of movies from pretraining, which can be used to answer these questions. To assess the impact of prior knowledge, we test different models using question-only inputs. The results are shown in the last two columns of Tab. 5. We find that ChatGPT-4o, Gemini-1.5-pro, and LongVA achieve high performance on ID questions with question-only inputs, with +22.5%, +14.0%, and +13.4% over random chance, respectively. This indicates that they can call upon prior knowledge to answer the questions. In contrast, Claude-3.5-sonnet and InternVL2-8B are less affected by prior knowledge. For almost all models tested, the performance with video and narration input is much higher than with question only, suggesting knowledge leakage alone is not sufficient for good performance on Causal2Needles.
@@ Appendix A Dataset Statistics
The videos in Causal2Needles are collected from the SyMoN (Sun et al., 2022) and YMS (Dogan et al., 2018) datasets, comprising movie recap videos and their corresponding narrations (98 movies from SyMoN and 94 from YMS). After integration, the videos in Causal2Needles have an average duration of 438 seconds, with 58.6 clips and 1,267.2 words per movie. The movie themes in our dataset are highly diverse. To quantify this diversity, we first use ChatGPT-4o to assign multi-labels to each narrative based on 11 predefined thematic categories, and then compute the frequency of each predicted label. The distribution is shown in Fig. 5 (a). As detailed in Sec. 3, we extract 747 cause-effect pairs from the narrations and construct 2-needle questions accordingly. We further analyze the distribution of their distances, defined as the number of clips between the cause and effect events, shown in Fig. 5 (b).
Appendix B Evaluation Details
@@
Current video understanding VLMs primarily rely on converting videos into sequences of images for processing. For example, LLaVA-Next-Video (Zhang et al., 2024b) samples 32 frames from a video, encodes them into visual tokens using a Vision Transformer (ViT) (Dosovitskiy et al., 2020), and feeds these tokens, along with textual instructions, into Vicuna-1.5 (Zheng et al., 2023) to generate responses. Interleaved VLMs (Chen et al., 2024; Li et al., 2024b) extend input formats to support text-image sequences.
Proprietary models such as Gemini-Pro (Team et al., 2024), GPT-4o (OpenAI, 2024a), and Claude-3.5 (Anthropic, 2024) benefit from substantially greater computational resources, enabling them to process millions of image tokens—achieving dense visual understanding at up to 2 frames per second. These models also support interleaved text-image inputs, offering stronger temporal and contextual reasoning capabilities over video content.
@@
Appendix C Additional Experiments
In this section, we provide a deeper analysis of several aspects: (1) the influence of the distance between cause and effect events; (2) textual grounding versus visual grounding; (3) the uniqueness of answers to Part 2 questions; and (4) non-stacked versus stacked image input settings.
@@ Appendix C Additional Experiments
[…]
@@
We conduct an additional experiment to verify if Part 2 questions have unique answers. During causal relationship extraction, an event graph may contain multiple nodes pointing to the same node, indicating that a single event can have multiple causes. This may introduce multiple ground truth answers to Part 2 questions, as they require retrieving the cause of an event.
Our experiment is designed as follows: among the $747$ causal relationships of 2-needle questions, we identify $99$ instances where the effect sentence is not uniquely paired with a single cause sentence. For each effect sentence in them, we feed the corresponding Part 2 question and cause sentences (on average, $2.02$ candidate causes per effect) to ChatGPT-4.1. Then, it is asked to select the cause sentence that contains the answer to the Part 2 question. To avoid bias, we randomize the order of the candidate causes for each query. The results show that out of the $99$ such cases, ChatGPT-4.1 selects the wrong cause sentence in only one instance. This indicates that the answer to Part 2 question is mostly unique.
@@
We explain the rationale for the design choice of inputting each video clip by stacking the video frames into a single image. We represent a video clip using five uniformly sampled frames. After that, we stack these frames from top to bottom as a single image as input to the VLMs. There are two reasons for the stacked approach. First, most open-source models have a limited window size and can only process 32 frames at a time. Therefore, without stacking, processing multiple video clips would quickly exceed the window size. Second, for proprietary models that can accept more frames as input, stacking the frames can help the model better capture and comprehend the story within a video clip. As verification, we conduct additional experiments on 2-needle VG questions using ChatGPT-4o and Gemini-1.5-pro with non-stacked frames as input. For a fair comparison, we instruct the model to output the frame number containing the answer and calculate the scene number by dividing the frame number by five, as every five frames belong to a single clip. The results are detailed in Tab. 5. We can observe that stacked frames improve performance significantly, indicating that the stacked method enhances visual understanding for VLMs.
Appendix D Limitations

### v3

@@ 4.4 Analysis of Dataset Bias
To prevent performance underestimation caused by the out-of-distribution VG question format, we introduce an image description (ID) format. However, an VLM may have memorized details of movies from pretraining, which can be used to answer these questions. To assess the impact of prior knowledge, we test different models using question-only inputs. The results are shown in the last two columns of Tab. 4. We find that ChatGPT-4o, Gemini-1.5-pro, and LongVA achieve high performance on ID questions with question-only inputs, with +22.5%, +14.0%, and +13.4% over random chance, respectively. This indicates that they can call upon prior knowledge to answer the questions. In contrast, Claude-3.5-sonnet and InternVL2-8B are less affected by prior knowledge. For almost all models tested, the performance with video and narration input is much higher than with question only, suggesting knowledge leakage alone is not sufficient for good performance. There remains substantial room for improvement (around 60–70%) in Causal2Needles, which could benefit from enhanced causal reasoning capability.
Acknowledgments
This research is supported by the RIE2025 Industry Alignment Fund – Industry Collaboration Projects (IAF-ICP) (Award I2301E0026), administered by A*STAR, by Alibaba Group and NTU Singapore through Alibaba-NTU Global e-Sustainability CorpLab (ANGEL), by the Nanyang Associate Professorship, and by the National Research Foundation Fellowship (NRFF13-2021-0006), Singapore.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: In the abstract, we clearly state our primary contribution, which is the design of a novel set of benchmark questions specifically created to evaluate models’ abilities in long-context video understanding. In the introduction, we highlight two key limitations of existing video understanding datasets to underscore the necessity and effectiveness of our proposed benchmark. We further summarize the performance of contemporary video large language models and identify their strengths and weaknesses. Thus, these claims represent the primary contributions of our paper, and these contributions are also summarized explicitly in the introduction.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: In Appendix Section E.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: This paper does not include theoretical results.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: For question generation, the procedure are introduced in Section 3. For model evaluation, We comprehensively introduced the models used in our experiments as well as the experimental procedures in Section 4. Additionally, we have publicly released both the code and benchmark dataset to facilitate reproduction and verification of our main experimental results.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: Following the guidelines, we have fully disclosed all information necessary to reproduce our main experimental results. We have publicly released both our benchmark dataset (available at: https://huggingface.co/datasets/causal2needles/Causal2Needles) and the code used for evaluating model performance (available at: https://github.com/jdsannchao/Causal2Needles). For proprietary models, we set the temperature parameter to 0 to minimize randomness in generation. For open-source models, we set do_sample to False to ensure deterministic outputs. These settings allow our reported results to be reliably reproduced.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: In Section 4.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [N/A]
  34. Justification: To minimize randomness and ensure reproducibility, all models were evaluated using fixed decoding settings. Specifically, we set temperature to 0 for proprietary models and do_sample to False for open-source models. As a result, model outputs are consistent across runs, and standard statistical significance tests or error bars are not applicable in this setting.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: In Appendix Section C.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: Yes, the research conducted in this paper fully conforms to the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: Our method primarily influences the evaluation of large video language models and may encourage the development of stronger video understanding capabilities. It does not have any direct societal impact, nor does it pose any foreseeable negative consequences.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: No additional safeguards were required, as the dataset is used strictly for its intended academic purpose of evaluating model performance. The datasets used in our study are all publicly available, originally sourced from YouTube videos. We do not introduce any new high-risk data or models. Our benchmark is constructed purely for academic evaluation purposes and does not contain sensitive, private, or potentially misusable content.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: All referenced datasets have been properly cited.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: All new assets introduced in the paper, including our benchmark dataset and evaluation scripts, are well documented.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: There is no crowd sourcing experiments.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [Yes]
  74. Justification: This paper does not involve any potential risks. For human evaluation of the dataset, we obtained IRB approvals.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [Yes]
  79. Justification: The use of LLMs is a central component of our methodology. LLMs are involved in every step of our question generation pipeline. We explicitly describe the role of LLMs in Section 3.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
@@ Appendix A Dataset Statistics
The videos in Causal2Needles are collected from the SyMoN (symon) and YMS (yms) datasets, comprising movie recap videos and their corresponding narrations (98 movies from SyMoN and 94 from YMS). After integration, the videos in Causal2Needles have an average duration of 438 seconds, with 58.6 clips and 1,267.2 words per movie. The movie themes in our dataset are highly diverse. To quantify this diversity, we first use ChatGPT-4o to assign multi-labels to each narrative based on 11 predefined thematic categories, and then compute the frequency of each predicted label. The distribution is shown in Fig. 6 (a). As detailed in Sec. 3, we extract 747 cause-effect pairs from the narrations and construct 2-needle questions accordingly. We further analyze the distribution of their distances, defined as the number of clips between the cause and effect events, shown in Fig. 6 (b).
Appendix B Event Graph Merging
To obtain comprehensive and long-range causal relationships, we merge the extracted local and global event graphs. The merging process simply takes the union of the causal relations from different event graphs. For example, suppose the output of Graph A is [(1, 5), (3, 10), (21, 27)]. Here, (1, 5) denotes a causal relation from event 1 to event 5. Further suppose the output of Graph B is [(1, 5), (3, 10), (6, 11)]. The merged result is the union [(1, 5), (3, 10), (6, 11), (21, 27)].
Appendix C Evaluation Details
@@
Current video understanding VLMs primarily rely on converting videos into sequences of images for processing. For example, LLaVA-Next-Video (zhang2024video_llavanext) samples 32 frames from a video, encodes them into visual tokens using a Vision Transformer (ViT) (dosovitskiy2020image_vit), and feeds these tokens, along with textual instructions, into Vicuna-1.5 (zheng2023judging_vicuna) to generate responses. Interleaved VLMs (chen2024far_internvl; li2024aria) extend input formats to support text-image sequences.
Proprietary models such as Gemini-Pro (team2024gemini), GPT-4o (openai2024chatgpt4o), and Claude-3.5 (anthropic2024claude3) benefit from substantially greater computational resources, enabling them to process millions of image tokens—achieving dense visual understanding at up to 2 frames per second. These models also support interleaved text-image inputs, offering stronger temporal and contextual reasoning capabilities over video content.
@@
Appendix D Additional Experiments
In this section, we provide a deeper analysis of several aspects: (1) the influence of the distance between cause and effect events; (2) textual grounding versus visual grounding; (3) the uniqueness of answers to Part 2 questions; (4) existence of counterfactual answers in ID 2-needle questions; and (5) non-stacked versus stacked image input settings.
@@ Appendix D Additional Experiments
First, the results show that the Part 2 question relies on joint understanding of both the cause and the effect sentences. We compute the conditional probability that GPT-4o answers SG Part 2 correctly given that it answers SG Part 1 correctly, $\mathrm{P}(\text{SG-Part2}={\text{correct}}\mid\text{SG-Part1}={\text{correct}})=82.1\%$. In contrast, the conditional probability that GPT-4o answers SG Part 2 correctly given that it answers SG Part 1 incorrectly, $\mathrm{P}(\text{SG-Part2}={\text{correct}}\mid\text{SG-Part1}={\text{incorrect}})=39.6\%$. The large gap between the two conditional probabilities demonstrates that performance on SG Part 2 significantly depends on SG Part 2. This suggests the existence of a sequential dependency between the two parts. The finding demonstrates that, at least in the textual modality, the two-needle questions benefit from the joint understanding of both needles.
Second, the dependence of VG performance on SG performance is limited. We compute the conditional probability that GPT-4o answers VG Part 2 correctly given that it answers SG Part 2 correctly, $\mathrm{P}(\text{VG-Part2 = \text{correct}}\mid\text{SG-Part2}={\text{correct}})=33.5\%$. In contrast, the conditional probability that the model answers VG Part 2 correctly given that it answers SG Part 2 incorrectly is $\mathrm{P}(\text{VG-Part2}={\text{correct}}\mid\text{SG-Part2}={\text{incorrect}})=25.4\%$. The difference between the two probabilities is small. This shows the correctness of SG Part 2 only makes minor contribution to VG Part 2. That is, even if the model is wrong on SG Part 2, it still has a decent chance of being correct on VG Part 2, relative to the alternative situation that the model is correct on SG Part 2. Hypothetically, a main reason for this phenomenon is that the model utilizes a separate mechanism for visual reasoning, which operates more or less independently of reasoning over the textual modality.
@@
We conduct an experiment to verify if Part 2 questions have unique answers. In causal relationship extraction, an event graph may contain multiple nodes pointing to the same node, indicating that a single event can have multiple causes. This may introduce multiple ground truth answers to Part 2 questions, as they require retrieving the cause of an event.
Our experiment is designed as follows: among the $747$ causal relationships of 2-needle questions, we identify $99$ instances where the effect sentence is not uniquely paired with a single cause sentence. For each effect sentence in them, we feed the corresponding Part 2 question and cause sentences (on average, $2.02$ candidate causes per effect) to ChatGPT-4.1. Then, it is asked to select the cause sentence that contains the answer to the Part 2 question. The results show that out of the $99$ such cases, ChatGPT-4.1 selects the wrong cause sentence in only one instance. This indicates that the answer to Part 2 question is mostly unique.
@@
Qwen2.5VL-7B non-stacked 2.1 11.0 0.0 13.8 4.4 0.4 8.0 7.7 0.2
stacked 27.6 17.7 5.0 11.2 18.9 1.9 19.4 18.3 3.4
Existence of Counterfactual Answers in ID 2-Needle Questions.
A counterfactual answer refers to an answer that is supported apparently by the input video and can be ruled out only by causal reasoning. If the video contains only the correct answer and no counterfactual answers, a neural network can easily rule out wrong answers without locating the required video clip.
For example, in Fig. 2, the image description question is “What tragedy was commemorated by the memorial service in Metropolis? When the tragedy that causes the memorial service occurs, what color were the energy blasts emanating from the creature? A. Yellow B. Blue C. Red D. White.” The correct answer is “Yellow”, which appears in the cause clip. “Blue” is a competing answer as a blue light is shown in another battle clip at 2’37 of the video. Due to the presence of both answers in the input video, to correctly answer the question, the model must identify the cause clip through causal reasoning.
We adopt two question design methods to ensure the existence of counterfactual answers in image description questions. First, we are deliberately ambiguous about the reference in the question wording, as discussed in Sec. 3.3. Second, we encourage the LLM to search for visually similar but potentially confusing content from adjacent clips, using the instructions in Appendix Fig. 14.
To demonstrate the existence of counterfactual answers, we conduct an experiment with the following setting: (1) We remove the cause clip from the input and any information that helps to locate the cause clip from the question, allowing the model to answer the question using other clips. For example, the aforementioned question becomes “ What color were the energy blasts emanating from the creature?”; (2) We replace the ground truth answer with “None is correct,” so if there are no counterfactual answers, the model should choose this option. In experiments with ChatGPT-4o, the model selects “None is correct” for only 17.8% of questions. This suggests that most image description questions in Causal2Needles have counterfactual answers.
@@
We explain the rationale for the design choice of inputting each video clip by stacking the video frames into a single image. We represent a video clip using five uniformly sampled frames. After that, we stack these frames from top to bottom as a single image as input to the VLMs. There are two reasons for the stacked approach. First, most open-source models have a limited window size and can only process 32 frames at a time. Therefore, without stacking, processing multiple video clips would quickly exceed the window size. Second, for proprietary models that can accept more frames as input, stacking the frames can help the model better capture and comprehend the story within a video clip. As verification, we conduct additional experiments on VG 2-needle questions using ChatGPT-4o, Gemini-1.5-pro and Qwen2.5VL-7B with non-stacked frames as input. For a fair comparison, we instruct the model to output the frame number containing the answer and calculate the clip number by dividing the frame number by five, as every five frames belong to a single clip. The results are detailed in Tab. 5. We can observe that stacked frames improve performance significantly, indicating that the stacked method enhances visual understanding.
Appendix E Limitations